### v2

@@ Abstract
Set-based state estimation plays a vital role in the safety verification of dynamical systems, which becomes significantly challenging when the system’s sensors are susceptible to cyber-attacks. Existing methods often impose limitations on the attacker’s capabilities, restricting the number of attacked sensors to be strictly less than half of the total number of sensors. This paper proposes a Secure Set-Based State Estimation (S3E) algorithm that addresses this limitation. The S3E algorithm guarantees that the true system state is contained within the estimated set, provided the initialization set encompasses the true initial state and the system is redundantly observable from the set of uncompromised sensors. The algorithm gives the estimated set as a collection of constrained zonotopes, which can be employed as robust certificates for verifying whether the system adheres to safety constraints. Furthermore, we demonstrate that the estimated set remains unaffected by attack signals of sufficiently large and also establish sufficient conditions for attack detection, identification, and filtering. This compels the attacker to inject only stealthy signals of small magnitude to evade detection, thus preserving the accuracy of the estimated set. When a few number of sensors (less than half) can be compromised, we prove that the estimated set remains bounded by a contracting set that converges to a ball whose radius is solely determined by the noise magnitude and is independent of the attack signals. To address the computational complexity of the algorithm, we offer several strategies for complexity-performance trade-offs. The efficacy of the proposed algorithm is illustrated through its application to a three-story building model.
@@ 1 Introduction
The interconnectedness of control systems makes them vulnerable to malicious attacks. Sensors are particularly susceptible to these attacks compared to other control system components. The compromised sensor measurements can significantly compromise system safety when employed for control.
To mitigate this risk and ensure accurate state estimation, several algorithms have been proposed recently. Known as resilient or secure state estimation, these algorithms leverage the redundancy of sensors to obtain estimates that converge to a neighborhood of the true state, even in the presence of additive sensor attacks. Secure state estimation algorithms (see, e.g., [14, 11, 19, 29, 16, 3, 9]) are designed to prevent corrupted sensor data from degrading estimation accuracy and maintain reasonable control performance. However, most methods proposed in the literature are point-based, i.e., the estimated state at any given time is a point in the state space. The estimation error bounds of point-based methods are not precise because of their dependence on comparison functions, which are known to be conservative. Moreover, the reliance of point-based methods on majority voting-type algorithms presents a fundamental limitation that only less than half the total number of sensors are allowed to be attacked at any given time. However, an adversary with ample resources would be capable of attacking more than half the number of sensors. While recent efforts investigating the incorporation of intermittent authentication protocols such as cryptography [18] alleviates this fundamental limit, its usage is a tradeoff between computation overhead and estimation accuracy.
1.1 Related Literature
To overcome the limitations of point-based methods and obtain tighter robustness guarantees, the set-based zonotopic filtering paradigm [1, 20, 33, 6, 5, 13, 2, 31] is promising. It has found many real-world applications, including fault diagnosis in industrial systems [7], underwater robotics [17], vehicle localization [8], and leakage detection in water distribution networks [32]. Moreover, in safety-critical applications, guaranteed state inclusion in a bounded set is crucial to verify whether the system adheres to the safety constraints. All these issues further motivate the need for set-based state estimation techniques under unknown system disturbances and measurement errors.
Different from zonotopic estimators, other set-based estimators include interval observers [30, 39], which estimate a box at every time instant, guaranteeing the true state’s inclusion. However, a box fails to capture the inter-dependencies between the state variables and is less accurate [32]. In contrast, zonotopic filters are shown to not only provide an accurate set-based estimate but are also computationally efficient.
Despite the growing importance of secure state estimation in the presence of adversarial attacks on sensors, the existing set-based methods have several shortcomings, including restrictive assumptions, limited robustness to stealthy attacks, and reliance on specific attack strategies. The reachability-based approach of [35] requires the full state vector to be measured by any subset of sensors with cardinality equal to the number of safe sensors. This is a highly restrictive assumption, and without it, the guarantees of estimation accuracy provided in [35] become excessively conservative. Recent works [26, 10, 22] employ standard interval-based or zonotopic filters by identifying and filtering the compromised sensors. However, this approach fails to exclude stealthy attacks on the sensors, which can easily evade the proposed attack detection procedure. As a result, stealthy attacks can significantly affect estimation accuracy by constantly injecting undetectable signals in the sensor measurements, which can accumulate over time and significantly degrade the estimation accuracy. Although this issue is addressed by [24, 23] and [37, 41, 25, 43], the proposed methods are either restricted to a particular attack strategy or rely on the assumption that the attack signal is bounded.
1.2 Our Contribution
In this paper, we propose a novel secure set-based state estimation (S3E) algorithm using zonotopic filtering that does not allow stealthy attacks to significantly corrupt the estimation performance. We also do not assume the boundedness of the attack signals or that the attacker resorts to a particular attack strategy. Moreover, to address the fundamental limitation of point-based secure estimators that strictly less than half the number of sensors can be attacked at any given time, we allow the attacker to compromise all but one sensors under the assumption that the system remains redundantly observable from the remaining uncompromised sensors. Subject to these assumptions, we show that the estimated set obtained from the S3E algorithm is guaranteed to contain the true state of the system.
Although the algorithm presented in this paper can also handle sensor faults, it was developed with adversarial attacks in mind. Attacks differ from faults in that an attack is an intelligently designed signal injected into the sensors to inflict a maximal negative impact on the system and/or evade detection, or in other words, is stealthy. For example, injecting small signals to corrupt the sensor measurements, which cannot be detected at any instant in time, but its effect can accumulate over a long time horizon. Another example is the so-called zero dynamics attack [34], where a specially designed sensor attack signal based on the unstable zero dynamics of the plant can grow unbounded while the plant’s state is driven away from the attack-free trajectory. Such an attack is directed close to the output null space such that the sensor measurement is close to zero (modulo noise). These intelligently designed attacks can cause conventional state estimation algorithms to provide inaccurate state estimates, where the estimation accuracy is dependent on the attack signals. We eliminate the undesirable effect of sensor attacks with the proposed S3E algorithm.
Our algorithm, presented in Section 4, operates in a series of steps at every time instant. Firstly, it calculates the time update by utilizing the model and the bounds on the process noise. Subsequently, subsets of time update consistent with the sensor measurements are computed through a measurement update. The measurement update sets correspond to the possible states in the time update that could have generated the obtained sensor measurements, provided that the exact realization of the measurement noise can be anywhere within the specified bounds described by a zonotope. Some attacked measurement update sets may turn out to be empty if the corresponding attack signal pushes the sensor measurement to a region that is inconsistent with the time update set. In practical scenarios, an attacker could compromise a large subset of sensors. To account for this, we create multiple agreement sets by intersecting various combinations of measurement update sets depending on the parameter of redundant observability of the system. By doing so, we can also eliminate measurement update sets that correspond to attacked sensors. We establish that at least one agreement set contains the true state of the system. Lastly, the estimated set is computed by taking the union of all the agreement sets.
The main strength of our algorithm is that it can handle attacks on different subsets of sensors at any given time. In addition, our approach can accommodate any cardinality of attacked sensors as long as they are less than the total number of sensors. The reason for this is that we do not rely on a simple majority vote among the sensors. Instead, we use a combination of intersections between multiple combinations of measurement update sets, which lead to the agreement sets. These combinations are compared with the model-based time update set in such a way that we can verify the validity of multiple agreement sets even if only one sensor is uncompromised. When at least half of the sensors can be compromised by the attacker, the accuracy of the point-based state estimation algorithms becomes highly susceptible. This is because, in this case, the attacker can easily manipulate the state estimate by injecting arbitrarily large signals. In contrast, our set-based algorithm is resilient to large attack signals as it can naturally detect and identify the corresponding compromised sensors through the measurement update sets or the agreement sets. In Section 5, we provide lower bounds on the magnitude of the attack signals that guarantee attack detection, identification, and/or filtering. Consequently, for the attacker to maintain both stealth and efficacy, they must resort to injecting attack signals of sufficiently small magnitudes.
A shortcoming of our algorithm is that its complexity may increase with time when stealthy attacks are executed, which in our paper are the attack signals with sufficiently small magnitude. Despite this, we argue in Section 5 that it is challenging for the attackers to achieve this worst-case scenario because it requires a complete understanding of both the system and the algorithm, as well as ample computational resources to compute a feasible attack within a single time step. Nonetheless, to tackle this complexity issue, we suggest various strategies to reduce the complexity and facilitate the implementation of the S3E algorithm.
Finally, for analyzing the stability and asymptotic behavior of the S3E algorithm, we integrate a secure point-based state estimator into our algorithm to prune the candidate estimated sets in Section 6. Because of employing a secure point-based estimator, this section relies on the assumption that strictly less than half of the sensors are compromised. However, this assumption is only needed when one requires that the radius of the estimated set is bounded by a contracting ball in a Euclidean space up to a margin that is dependent only on the disturbance and measurement noise. If this assumption is violated, then we can only ensure the boundedness of the radius of the estimated set at every time instant.
In summary, our contributions in this paper include a novel secure set-based state estimation algorithm [^1] The code is available online in our GitHub repository https://github.com/aalanwar/Secure-Set-Based-Estimation. that guarantees the following:
When up to $p-1$ out of total $p$ sensors can be compromised:
  1. 1.
    guaranteed state inclusion, i.e., the true state is always guaranteed to be inside the estimated set;
  2. 2.
    secure state estimation, i.e., large attack signals are automatically discarded, while the impact of stealthy attacks on the estimation accuracy is bounded;
  3. 3.
    sufficient conditions for detection, identification, and filtering of attacks on sensors;
  4. 4.
    methods to address the algorithm’s complexity with minimal compromise on the estimation accuracy.
When strictly less than $p/2$ sensors can be compromised:
  1. (5)
    stability guarantees, i.e., the estimated set remains bounded by a contracting ball whose asymptotic radius only depends on bounds of the process and measurement noise.
To this end, we assume that the system is redundantly observable, i.e., observable from every possible combination of sensors with cardinality less than or equal to the number of safe sensors. This assumption is more realistic than the one in our previous work [28], which required observability from every sensor. Also, the analysis under this assumption turns out to be significantly different and non-trivial compared to [28]. Furthermore, we provide stability guarantees and an attack detection and identification algorithm, which is missing in [28]. Finally, we demonstrate the effectiveness of our proposed method through an illustrative example and a more practical example of a three-story building structure [38] during an earthquake and when an adversary has compromised multiple sensors.
1.3 Outline
@@
The rest of the paper is organized as follows. Section 2 defines the notations and provides the required preliminaries on set representations used in this paper. The main assumptions and the problem are stated in Section 3. Section 4 presents the secure zonotopic state estimation algorithm, provides the inclusion guarantees, and discusses methods to reduce the algorithm’s complexity. Section 5 provides lower bounds on the attacks that can be detected, identified, and/or filtered. Section 6 provides the stability analysis of the estimation algorithm. Finally, Section 7 demonstrates the effectiveness of the proposed algorithm through simulation examples, and Section 8 concludes the paper.
2 Preliminaries
This section establishes the notation used consistently throughout the paper. Additionally, it briefly introduces zonotopes, constrained zonotopes, and their relevant properties and operations employed in the proposed algorithm.
2.1 Notations
@@
The set of real numbers and integers are denoted by $\mathbb{R}$ and $\mathbb{Z}$, respectively. We let $\mathbb{Z}_{\geq i}\coloneq\{i,i+1,i+2,\dots\}$ and $\mathbb{Z}_{[i,j]}\coloneq\{i,i+1,i+2,\dots,j\}$ for $j\geq i$. The Euclidean and maximum norms of a vector $x\in\mathbb{R}^{n}$ are denoted as $\|x\|=\sqrt{x^{\top}x}$ and $\|x\|_{\infty}\coloneq\max_{i\in\{1,\dots,n\}}|x_{i}|$. The vectors of zeros and ones of dimensions $n\times 1$ are denoted as $0_{n}$ and $1_{n}$, respectively. Similarly, $0_{n\times k}$ and $1_{n\times k}$ denote matrices of zeros and ones of dimensions $n\times k$. The identity matrix is $I_{n}\in\mathbb{R}^{n\times n}$. Given a signal $v:\mathbb{Z}_{\geq 0}\rightarrow\mathbb{R}^{n}$, we denote its restriction to the interval $[0,k]$ by $v_{[0,k]}$, for some $k\in\mathbb{Z}_{\geq 0}$. For a finite set $\mathcal{S}$, $|\mathcal{S}|$ denotes its cardinality. The Cartesian product between two compact sets $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ is given by $\mathcal{S}_{1}\times\mathcal{S}_{2}$. The notation (${n\choose k}=\frac{n!}{n!(n-k)!}$ is the number of possible combinations of $k$ out of $n$ elements.
2.2 Set Representations
In this subsection, we introduce the set representations relevant in this paper. Particularly, we define zonotopes and constrained zonotopes, and some set-based operations along with their relevant properties.
2.2.1 Zonotope
@@
Given a center $c_{z}\in\mathbb{R}^{n}$ and a generator $G_{z}\in\mathbb{R}^{n\times\xi_{z}}$, a zonotope $\mathcal{Z}\subset\mathbb{R}^{n}$ is a set
@@
where $\xi_{z}$ is the number of generators of $\mathcal{Z}$. That is, a zonotope is an affine transformation of a unit hypercube $\mathcal{H}(0_{n},1)=[-1,1]^{\xi_{z}}$ centered at $0_{n}$ and with radius $1$, where $\xi_{z}\in\mathbb{Z}_{\geq 1}$ is the dimension of the hypercube. Since a zonotope can be completely characterized by its center and generator, the notation $\mathcal{Z}=\langle c_{z},G_{z}\rangle$ is used throughout the paper for brevity.
2.2.2 Basic operations on zonotopes
A matrix $L\in\mathbb{R}^{n^{\prime}\times n}$ multiplied with a zonotope $\mathcal{Z}$ yields a linearly transformed zonotope $L\mathcal{Z}=\langle Lc_{z},LG_{z}\rangle$. Given two zonotopes $\mathcal{Z}_{1}=\langle c_{z_{1}},G_{z_{1}}\rangle$ and $\mathcal{Z}_{2}=\langle c_{z_{2}},G_{z_{2}}\rangle$, each being a subset of $\mathbb{R}^{n}$, their Minkowski sum is given by
@@ 2.2.2 Basic operations on zonotopes
Similarly, the Cartesian product of two zonotopes is defined and computed as
$\mathcal{Z}_{1}\times\mathcal{Z}_{2}$ :-$\coloneq\left\{{\left[{\begin{array}[]{c}{{z_{1}}}\\ {{z_{2}}}\end{array}}\right]}:{z_{1}}\in{\mathcal{Z}_{1}},{z_{2}}\in{\mathcal{Z}_{2}}\right\}:$
$=\left\langle\left[\begin{array}[]{c}c_{z_{1}}\\ c_{z_{2}}\end{array}\right],\left[\begin{array}[]{cc}G_{z_{1}}&0\\ 0&G_{z_{2}}\end{array}\right]\right\rangle.$
2.2.3 Constrained zonotope
A constrained zonotope is a set
@@ 2.2.3 Constrained zonotope
where $c_{z}\in\mathbb{R}^{n}$, $G_{z}\in\mathbb{R}^{n\times\xi_{z}}$, $A_{z}\in\mathbb{R}^{m\times\xi_{z}}$ and $b_{z}\in\mathbb{R}^{m}$ with $m,n\in\mathbb{Z}_{\geq 1}$. That is, a constrained zonotope is an affine transformation of the linearly constrained unit hypercube $\{\beta_{z}\in[-1,1]^{\xi_{z}}:A_{z}\beta_{z}=b_{z}\}$. We use the notation $\mathcal{Z}=\langle c_{z},G_{z},A_{z},b_{z}\rangle$ for a constrained zonotope.
2.2.4 Generalized intersection
@@
Let $\mathcal{Z}=\langle c_{z},G_{z},A_{z},b_{z}\rangle$ with $c_{z}\in\mathbb{R}^{n_{z}}$, $G_{z}\in\mathbb{R}^{n_{z}\times\xi_{z}}$, $A_{z}\in\mathbb{R}^{m_{z}\times\xi_{z}}$, and $b_{z}\in\mathbb{R}^{m_{z}}$, and let $\mathcal{Y}=\langle c_{y},G_{y},A_{y},b_{y}\rangle\subset\mathbb{R}^{n_{y}}$ with $c_{y}\in\mathbb{R}^{n_{y}}$, $G_{y}\in\mathbb{R}^{n_{y}\times\xi_{y}}$, $A_{y}\in\mathbb{R}^{m_{y}\times\xi_{y}}$, and $b_{y}\in\mathbb{R}^{m_{y}}$, then [33] provides the following expression for computing the generalized intersection:
$\mathcal{Z}\cap_{M}\mathcal{Y}=\left\langle c_{z},\left[\begin{array}[]{cc}G_{z}&0_{n_{z}\times\xi_{y}}\end{array}\right],\left[\begin{array}[]{cc}A_{z}&0_{m_{z}\times\xi_{y}}\\ 0_{m_{y}\times\xi_{z}}&A_{y}\\ MG_{z}&-G_{y}\end{array}\right],\left[\begin{array}[]{c}b_{z}\\ b_{y}\\ c_{y}-Mc_{z}\end{array}\right]\right\rangle.$ (2)
2.2.5 Radius of a zonotope
The radius of a zonotope, or a constrained zonotope, $\mathcal{Z}\subset\mathbb{R}^{n}$ is computed as follows
$\mathrm{rad}(\mathcal{Z})=\min\Delta~{}\text{subject to}~{}\mathcal{Z}\subseteq\mathcal{H}(c_{z},\Delta)$
where $\Delta$ is the radius of a minimal hypercube of dimension $n$ denoted by
$\mathcal{H}(c_{z},\Delta)\coloneq\langle c_{z},\Delta I_{n}\rangle$
which is centered at $c_{z}$ and inscribes $\mathcal{Z}$. Notice that for any point $p\in\mathcal{Z}$, it holds that

### v3

@@ Abstract
Set-based state estimation provides guaranteed state inclusion certificates that are crucial for the safety verification of dynamical systems. However, when system sensors are subject to cyberattacks, maintaining both safety and security guarantees becomes a fundamental challenge that existing point-based secure state estimation methods cannot adequately address due to their inherent inability to provide state inclusion certificates. This paper introduces a novel approach that simultaneously ensures safety guarantees through guaranteed state inclusion and security guarantees against sensor attacks, without imposing conservative restrictions on system operation. We propose a Secure Set-based State Estimation (S3E) algorithm that maintains the true system state within the estimated set under sensor attacks, provided the initialization set contains the initial state and the system remains observable from the uncompromised sensor subset. The algorithm gives the estimated set as a collection of constrained zonotopes (agreement sets), which can be employed as robust certificates for verifying whether the system adheres to safety constraints. Furthermore, we demonstrate that the estimated set remains unaffected by attack signals of sufficiently large magnitude and also establish sufficient conditions for attack detection, identification, and filtering. This compels the attacker to only inject signals of small magnitudes to evade detection, thus preserving the accuracy of the estimated set. To address the computational complexity of the algorithm, we offer several strategies for complexity-performance trade-offs. The efficacy of the proposed algorithm is illustrated through several examples, including its application to a three-story building model.
@@ 1 Introduction
Secure state estimation is critical for the reliable operation of cyber-physical systems, where control performance and safety depend directly on accurate state knowledge. Adversarial manipulation of sensor networks through false data injection can compromise system stability and lead to operational failures [35]. This vulnerability presents significant challenges for critical infrastructure systems, including power grids, water distribution networks, autonomous vehicles, and industrial automation [18].
Modern control systems exhibit inherent vulnerabilities due to their distributed architectures and dependence on networked communication infrastructure. These characteristics enable adversaries to manipulate sensor measurements through false data injection attacks, creating divergence between reported and actual system states. Research demonstrates that sophisticated attacks can maintain stealth characteristics while progressively degrading system performance [37, 42]. The evolution of such attack strategies, coupled with increased system connectivity, drives the need for resilient estimation algorithms capable of maintaining state accuracy under sensor compromise.
Traditional approaches to secure state estimation have focused on point-based methods that produce single estimates of the system state at each time instant. These algorithms typically employ redundant sensors and leverage techniques such as majority voting or optimization-based outlier detection to identify and exclude corrupted measurements [14, 11, 20, 31, 16, 3, 9]. While these methods have shown promise in specific scenarios, they suffer from fundamental limitations that restrict their applicability. Most notably, point-based secure estimators require that strictly fewer than half of the sensors remain uncompromised. This constraint may be violated by adversaries with sufficient resources to coordinate large-scale attacks. Furthermore, these methods produce conservative error bounds that limit their practical applicability.
Recent efforts to address these limitations have explored the integration of cryptographic authentication mechanisms to periodically verify the integrity of sensor data [23, 19]. However, the computational overhead associated with cryptographic operations creates challenging tradeoffs between security guarantees and the real-time performance requirements of control systems. This overhead becomes particularly problematic in resource-constrained embedded systems where computational capacity must be carefully allocated between control tasks and security operations.
1.1 Set-based Estimation Framework
Beyond merely detecting attacks or providing state estimates, safety-critical systems require formal verification that the system remains within safe operating bounds even under sensor attacks. Point-based methods fail to address this verification challenge as they cannot represent the complete space of feasible system states given measurement data and attack constraints without excessive conservatism. This limitation becomes critical in industrial processes operating with minimal safety margins, where minor estimation uncertainties may result in constraint violations.
Set-based estimation techniques, particularly those employing zonotopic representations, offer a fundamentally different approach to secure state estimation that addresses many limitations of point-based methods while naturally supporting safety verification [1, 21, 34, 6, 5, 13, 2, 32]. These methods compute bounded sets containing all feasible system states rather than single estimates, yielding inclusion guarantees necessary for constraint verification. Moreover, set-based estimation methods naturally embed uncertainties and noise in cases where their bounds are known. The zonotopic representation balances computational efficiency with tight uncertainty bounds, making real-time implementation feasible.
The key advantage of set-based methods for safety verification lies in their ability to propagate uncertainty through system dynamics while maintaining guaranteed bounds. These methods compute reachable sets encompassing all trajectories consistent with measurements and bounded disturbances, enabling verification of safety constraint satisfaction. This capability proves essential against stealthy attacks that evade immediate detection while gradually compromising system safety. Applications spanning fault diagnosis in industrial systems [7], underwater robotics [17], vehicle localization [8], and water network leak detection [33] demonstrate the practical value of guaranteed uncertainty bounds for operational safety and performance verification.
1.2 Current Limitations in Secure Set-Based Estimation
Despite their theoretical advantages for safety verification, existing set-based methods for secure state estimation face practical challenges. The reachability-based approach of [36] imposes that any subset of uncompromised sensors can measure the full state vector. Without such strong observability conditions, the estimation bounds become excessively conservative, potentially encompassing regions too large for effective control or meaningful safety verification. Similarly, our previous work [30] on secure state estimation considered restrictive assumptions requiring observability from every individual sensor and was limited to systems that were bounded input-bounded state stable, which significantly constrains the class of systems that can be analyzed.
Alternative approaches that attempt to identify and filter compromised sensors before applying standard zonotopic filters [28, 10, 24] fail to address stealthy attacks. These attacks inject strategically designed signals that remain within expected bounds at each time step while causing significant cumulative deviation in state estimates. Although recent work has begun to address stealthy attacks [26, 25, 39, 43, 27, 45], existing methods either restrict their analysis to specific attack strategies or assume known bounds on attack signal magnitudes. Sophisticated adversaries can readily violate these assumptions.
Furthermore, the fundamental assumption of requiring a majority of uncompromised sensors remains unaddressed in most set-based approaches. This assumption becomes increasingly problematic as systems scale and adversaries gain greater capabilities to compromise distributed sensor networks. These limitations necessitate novel algorithms that leverage the geometric properties of set-based representations to maintain state estimation accuracy and safety guarantees under extensive sensor compromise.
1.3 Contributions of This Work
This paper presents a Secure Set-Based State Estimation (S3E) algorithm that provides guaranteed state inclusion under adversarial sensor attacks while fundamentally relaxing the assumptions required by existing methods. Our approach leverages the redundancy inherent in zonotopic representations to maintain accurate state estimates and verify safety constraints even when all but one sensor can be compromised, requiring only that the system remains observable through any combination of remaining safe sensors.
The S3E algorithm executes four sequential operations per time step. The time update computes feasible states based on system dynamics and bounded process noise. Measurement updates generate state sets consistent with sensor observations across different sensor subsets. Agreement sets form through strategic intersections of measurement updates, leveraging redundancy to exclude compromised sensor data. The final estimated set comprises the union of valid agreement sets, ensuring true state inclusion and enabling safety verification.
Our theoretical analysis establishes several key properties of the S3E algorithm with direct implications for safety verification. We derive explicit lower bounds on attack signal magnitudes that ensure detection at different stages of the estimation process, demonstrating that adversaries face fundamental constraints in designing attacks that compromise safety without detection. We characterize the conditions under which attacked sensors can be identified and filtered, providing system designers with quantitative metrics for evaluating security and safety margins. Additionally, we develop computational methods to verify estimated set containment within safety constraints, supporting real-time monitoring under active attacks.
1.4 Outline of the Paper
@@
The rest of the paper is organized as follows. Section 2 provides the required preliminaries and Section 3 states the main assumptions and the problem. Section 4 presents the S3E algorithm, provides the inclusion guarantees, and discusses methods to reduce the algorithm’s complexity. Section 5 provides lower bounds on the attacks that can be detected, identified, and/or filtered. Section 6 demonstrates the proposed algorithm through simulation examples. Finally, Section 7 provides the concluding remarks.
2 Notation and Preliminaries
@@
The set of real numbers and integers are denoted by $\mathbb{R}$ and $\mathbb{Z}$, respectively. We let $\mathbb{Z}_{\geq i}\coloneq\{i,i+1,i+2,\dots\}$ and $\mathbb{Z}_{[i,j]}\coloneq\{i,i+1,i+2,\dots,j\}$ for $j\geq i$. The Euclidean and maximum norms of a vector $x\in\mathbb{R}^{n}$ are denoted as $\|x\|$ and $\|x\|_{\infty}$, respectively. The vector and matrix of zeros are denoted as $0_{n}\in\mathbb{R}^{n}$ and $0_{n\times k}\in\mathbb{R}^{n\times k}$. The identity matrix is $I_{n}\in\mathbb{R}^{n\times n}$. For a finite set $\mathcal{S}$, $|\mathcal{S}|$ denotes its cardinality. The Cartesian product between two sets $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ is denoted by $\mathcal{S}_{1}\times\mathcal{S}_{2}$.
@@
Given a center $c_{z}\in\mathbb{R}^{n}$ and a generator $G_{z}\in\mathbb{R}^{n\times\xi_{z}}$, a zonotope $\mathcal{Z}=\langle c_{z},G_{z}\rangle$ is a set
@@
where $\xi_{z}$ is the number of generators of $\mathcal{Z}$. That is, a zonotope is an affine transformation of a unit hypercube $\mathcal{H}(0_{n},1)=[-1,1]^{\xi_{z}}$ centered at $0_{n}$ and with radius $1$, where $\xi_{z}\in\mathbb{Z}_{\geq 1}$ is the dimension of the hypercube.
A matrix $L\in\mathbb{R}^{n^{\prime}\times n}$ multiplied with a zonotope $\mathcal{Z}$ yields a linearly transformed zonotope $L\mathcal{Z}=\langle Lc_{z},LG_{z}\rangle$. Given two zonotopes $\mathcal{Z}_{1}=\langle c_{z_{1}},G_{z_{1}}\rangle$ and $\mathcal{Z}_{2}=\langle c_{z_{2}},G_{z_{2}}\rangle$ in $\mathbb{R}^{n}$, their Minkowski sum is given by
@@
The Cartesian product of two zonotopes is given by
$\mathcal{Z}_{1}\times\mathcal{Z}_{2}$ $\coloneq\left\{{\left[{\begin{array}[]{c}{{z_{1}}}\\ {{z_{2}}}\end{array}}\right]}:{z_{1}}\in{\mathcal{Z}_{1}},{z_{2}}\in{\mathcal{Z}_{2}}\right\}=\left\langle\left[\begin{array}[]{c}c_{z_{1}}\\ c_{z_{2}}\end{array}\right],\left[\begin{array}[]{cc}G_{z_{1}}&0\\ 0&G_{z_{2}}\end{array}\right]\right\rangle.$
A constrained zonotope $\mathcal{Z}=\langle c_{z},G_{z},A_{z},b_{z}\rangle$ is a set
@@
where $c_{z}\in\mathbb{R}^{n}$, $G_{z}\in\mathbb{R}^{n\times\xi_{z}}$, $A_{z}\in\mathbb{R}^{m\times\xi_{z}}$ and $b_{z}\in\mathbb{R}^{m}$ with $m,n\in\mathbb{Z}_{\geq 1}$. That is, a constrained zonotope is an affine transformation of the linearly constrained unit hypercube $\{\beta_{z}\in[-1,1]^{\xi_{z}}:A_{z}\beta_{z}=b_{z}\}$.
@@
Let $\mathcal{Z}=\langle c_{z},G_{z},A_{z},b_{z}\rangle\subset\mathbb{R}^{n_{z}}$ with $c_{z}\in\mathbb{R}^{n_{z}}$, $G_{z}\in\mathbb{R}^{n_{z}\times\xi_{z}}$, $A_{z}\in\mathbb{R}^{m_{z}\times\xi_{z}}$, and $b_{z}\in\mathbb{R}^{m_{z}}$, and let $\mathcal{Y}=\langle c_{y},G_{y},A_{y},b_{y}\rangle\subset\mathbb{R}^{n_{y}}$ with $c_{y}\in\mathbb{R}^{n_{y}}$, $G_{y}\in\mathbb{R}^{n_{y}\times\xi_{y}}$, $A_{y}\in\mathbb{R}^{m_{y}\times\xi_{y}}$, and $b_{y}\in\mathbb{R}^{m_{y}}$, then the generalized intersection can be computed as, see [34],
$\mathcal{Z}\cap_{M}\mathcal{Y}=\Big{\langle}c_{z},\left[\begin{array}[]{cc}G_{z}&0_{n_{z}\times\xi_{y}}\end{array}\right],\left[\begin{array}[]{cc}A_{z}&0_{m_{z}\times\xi_{y}}\\ 0_{m_{y}\times\xi_{z}}&A_{y}\\ MG_{z}&-G_{y}\end{array}\right],\left[\begin{array}[]{c}b_{z}\\ b_{y}\\ c_{y}-Mc_{z}\end{array}\right]\Big{\rangle}.$ (2)
The radius of a zonotope, or a constrained zonotope, $\mathcal{Z}\subset\mathbb{R}^{n}$ is defined as
$\mathrm{rad}(\mathcal{Z})\coloneq\min\Delta~\text{subject to}~\mathcal{Z}\subseteq\mathcal{H}(c_{z},\Delta)$
where $\Delta$ is the radius of a minimal hypercube of dimension $n$ denoted by $\mathcal{H}(c_{z},\Delta)\coloneq\langle c_{z},\Delta I_{n}\rangle$, which is centered at $c_{z}$ and inscribes $\mathcal{Z}$. Notice that for any point $p\in\mathcal{Z}$, it holds that